# Rethinking Pruning Large Language Models: Benefits and Pitfalls of Reconstruction Error Minimization

### v1

@@ Abstract
This work suggests fundamentally rethinking the current practice of pruning large language models (LLMs). The way it is done is by divide and conquer: split the model into submodels, sequentially prune them, and reconstruct predictions of the dense counterparts on small calibration data one at a time; the final model is obtained simply by putting the resulting sparse submodels together. While this approach enables pruning under memory constraints, it generates high reconstruction errors. In this work, we first present an array of reconstruction techniques that can significantly reduce this error by more than $90\%$. Unwittingly, however, we discover that minimizing reconstruction error is not always ideal and can overfit the given calibration data, resulting in rather increased language perplexity and poor performance at downstream tasks. We find out that a strategy of self-generating calibration data can mitigate this trade-off between reconstruction and generalization, suggesting new directions in the presence of both benefits and pitfalls of reconstruction for pruning LLMs.
@@
Sungbin Shin1thanks: Work partly done as a student researcher at Google  Wonpyo Park2  Jaeho Lee1,2  Namhoon Lee1,2 1POSTECH  2Google {ssbin4,jaeho.lee,namhoonlee}@postech.ac.kr wppark@google.com
@@ 1 Overview
Figure 1: (a) Reconstruction techniques significantly reduce the compounding errors and lead to a substantial reduction of error in the final block. (b) Minimizing reconstruction error may not always be ideal since models can overfit calibration data. Self-generated data effectively mitigates this issue and decreases test error, perplexity, and error rates for downstream tasks (here, the lower values indicate better results).
@@
In this work, we address the reconstruction error minimization for pruning LLMs with the following three major pillars. First, we focus on developing various engineering techniques to reduce this error. These are inspired to lessen the suboptimality of subsolutions by incorporating different levels of extension schemes. Second, we suggest that reducing this error is not necessarily favorable, however. Our extensive experimental results indicate that it is possibly due to overfitting, given limited calibration data and high problem complexity. Third, we present useful strategies to potentially mitigate the risk of reconstruction and improve generalization. This is based on the self-generation of calibration data.
Briefly, this work investigates the benefits and pitfalls of the reconstruction error minimization scheme for pruning LLMs. To our best knowledge, this trade-off has not been explicitly identified or studied before, suggesting rethinking the current practice. Our initial investigations may shed light on some potential future research directions. We summarize our results in Figure 1.
@@ 2 Reconstruction techniques
The seminal work of Frantar and Alistarh (2023) proposes to reconstruct predictions layer-wise based on least squares. By removing non-linearity this approach yields a closed-form solution, yet it can create a high reconstruction error since the system is highly underdetermined (i.e., there are much more parameters than calibration data). To reduce compounding errors, we first consider extending the unit of optimization target from a layer to a block of layers. Specifically, this means a block-wise reconstruction (br) which can be formulated as follows:
@@
where $g_{i}$ refers to the $i$-th block of layers (e.g., a Transformer block) in which we have the optimization variables $w_{i}$, and $x_{i}$ denotes the inputs which originally come from calibration data; here, the pruning mask $\bar{m}$ is fixed assuming that it is already obtained from an arbitrary pruning method. I.e., the goal is to update variables in each block to minimize the extended reconstruction errors. We solve this problem iteratively using the standard gradient-based method. We note that a similar approach is also proposed in the concurrent work of Guo et al. (2024), and we find that br is extremely effective in reducing reconstruction errors in Section 3.1. We illustrate the idea of br in Figure 2.
@@
Another way we consider to further reduce reconstruction errors is to extend the reconstruction unit from a block to multiple blocks and stitch the solutions in between by connecting via the adjacent block. Specifically, this means that now $g$ in (2) becomes a composite of multiple blocks, say $h$, and we ensure $h$ overlaps; more precisely, $h_{i}=g_{i}\circ g_{i-1}$ and $h_{i+1}=g_{i+1}\circ g_{i}$ for two blocks, and so on. This way, namely cross-block reconstruction or cr (Ding et al., 2023), we can potentially bridge between subsolutions by taking into account some interaction between adjacent blocks, and hence, reduce the compounding errors. We illustrate the idea of cr in Figure 2. However, albeit effectively reducing the error, we find that this method can often lead to overfitting. We discuss this in detail in Section 3.2.
@@ 3.1 Reconstruction error
We first evaluate the effectiveness of the reconstruction techniques in reducing the error. Here, we focus on pruning LLaMA-7B (Touvron et al., 2023) and OPT-125M (Zhang et al., 2022) to unstructured $50\%$ sparsity with three pruning methods: SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024), and Magnitude (Han et al., 2015). For each pruning method, we examine four reconstruction strategies: layer-wise reconstruction (lr), block-wise reconstruction (br), block-wise reconstruction with global propagation (br+gp), and cross-block reconstruction with global propagation (br+gp+cr). Following convention, we use $256$ calibration data randomly sampled from C4 (Raffel et al., 2020) each containing $1024$ tokens. We run the Adam optimizer for $10$ epochs (see Appendix A for details). The results are presented in Figure 3.
@@ 3.2 Generalization performance
We first find that the perplexity effectively decreases with br and gp; the value reduces across all test cases including different models, pruning methods, and datasets. Unexpectedly, however, the perplexity rather increases when we add cr despite the reduced reconstruction error. We also observe a similar trend in zero-shot performance for Wanda and Magnitude pruning, with mean accuracy increasing by a large margin with br and gp but decreasing with cr. Interestingly, for SparseGPT, reconstruction techniques do not generally help zero-shot performance. We hypothesize that it is because SparseGPT already conducts fairly heavy optimization compared to other methods, and applying further reconstruction on particular calibration data may not help improve zero-shot performance since it is more sensitive to distribution shift. Furthermore, we find that such overfiting tends to occur more for LLaMA-7B than OPT-125M (see Table 2). This is possibly due to model size; i.e., given the same amount of (limited) calibration data, over-optimizing can make large models more likely to overfit and lead to poor generalization.
@@ 4 Further Exploration
We have seen that reconstruction techniques can lead to undesirable overfitting. Here we explore potential ways to alleviate this risk. In particular, we identify that the calibration data is highly limited in two aspects: it is too little (compared to optimization variables) and does not represent the training data (as it is arbitrarily given); the former is related to the general representation-generalization complexity trade-off, and the latter is about whether the reconstruction can mimic the behavior of the original model. To this end, we reflect on the fact that what we are dealing with is a generative (language) model, meaning that we can create calibration data that is potentially much bigger in size and closer to the original distribution. We find that this self-generation technique has recently been proposed in other contexts (Meng et al., 2022; Ye et al., 2022; Liu et al., 2023), and thus, follow the process therein to produce high-quality text data. Using that, we perform reconstruction again, and the results are reported in Figure 4. We observe that making use of more self-generated calibration data (without unfairly violating the given setting) reduces both test error and perplexity, mitigating overfitting quite effectively.
@@ 6 Limitations
There remain several limitations in our experiments and we plan to address these in future work. First, our main experiments are limited to LLaMA-7B and OPT-125M. We will scale up our experiments to much larger models of up to 70B parameters and different architectures including Mixtral (Jiang et al., 2024) or Gemma (Team et al., 2024). Next, reconstruction techniques br, gp, and cr require additional memory compared to lr, although they still use much less memory compared to model-level reconstruction of solving (1). We plan to introduce parameter-efficient optimization (Hu et al., 2022) to alleviate this increased memory burden. Finally, although the self-generation of calibration data effectively mitigates overfitting, it requires more computation for reconstruction. We also find that some portions of the generated texts are far from plain English texts and thus may not serve as good calibration data (see Table 5 of Appendix C for the examples). In this regard, we believe that reducing the number of these irrelevant examples and generating only a few number of high-quality texts can be a potential way to improve performance and increase efficiency.
7 Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)) and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2023-00210466, RS-2023-00265444, RS2023-0021371). Sungbin Shin was supported by Kwanjeong Educational Foundation Scholarship.

### v2

@@ Abstract
This work suggests fundamentally rethinking the current practice of pruning large language models (LLMs). The way it is done is by divide and conquer: split the model into submodels, sequentially prune them, and reconstruct predictions of the dense counterparts on small calibration data one at a time; the final model is obtained simply by putting the resulting sparse submodels together. While this approach enables pruning under memory constraints, it generates high reconstruction errors. In this work, we first present an array of reconstruction techniques that can significantly reduce this error by more than $90\%$. Unwittingly, however, we discover that minimizing reconstruction error is not always ideal and can overfit the given calibration data, resulting in rather increased language perplexity and poor performance at downstream tasks. We find out that a strategy of self-generating calibration data can mitigate this trade-off between reconstruction and generalization, suggesting new directions in the presence of both benefits and pitfalls of reconstruction for pruning LLMs. [^1] Our code is available at https://github.com/LOG-postech/rethinking-LLM-pruning.
@@
Sungbin Shin1thanks: Work partly done as a student researcher at Google  Wonpyo Park2  Jaeho Lee1,2,3  Namhoon Lee1,2,3 1POSTECH  2Google  3Yonsei University {ssbin4,jaeho.lee,namhoonlee}@postech.ac.kr wppark@google.com
@@ 1 Refer to caption
Figure 1: (a) Reconstruction techniques significantly reduce the compounding errors and lead to a substantial reduction of error in the final block. Reconstruction o and x refer to the results with and without the proposed reconstruction techniques (br, gp, cr) respectively. (b) Minimizing reconstruction error may not always be ideal since models can overfit calibration data (we show this in Section 3.2). Using our self-generated calibration data in the reconstruction process mitigates this issue quite effectively by decreasing test error, perplexity, and error rates for downstream tasks.
@@
In this work, we address the reconstruction error minimization for pruning LLMs with the following three major pillars. First, we focus on developing various engineering techniques to reduce this error. These are inspired to lessen the suboptimality of subsolutions by incorporating different levels of extension schemes. Second, we suggest that reducing this error is not necessarily favorable, however. Our extensive experimental results indicate that it is possibly due to overfitting, given limited calibration data and high problem complexity. Third, we present useful strategies to potentially mitigate the risk of reconstruction and improve generalization. This is based on what we call the self-generation of calibration data.
Briefly, this work investigates the benefits and pitfalls of the reconstruction error minimization scheme for pruning LLMs. To our best knowledge, this trade-off has not been explicitly identified or studied before, thereby suggesting rethinking the current practice. Our initial investigations may shed light on some potential future research directions. We summarize our main results in Figure 1.
@@ 2 Reconstruction Techniques
The seminal work of Frantar and Alistarh (2023) proposes to reconstruct predictions layer-wise based on least squares. By removing non-linearity this approach yields a closed-form solution. However, we find that this can create a high reconstruction error since the system is highly underdetermined (i.e., there are much more parameters than calibration data). To reduce compounding errors, we first consider extending the unit of optimization target from a layer to a block of layers. Specifically, this means a block-wise reconstruction (br) which can be formulated as follows:
@@
where $g_{i}$ refers to the $i$-th block of layers (e.g., a Transformer block) in which we have the optimization variables $w_{i}$, and $x_{i}$ denotes the inputs to the $i$-th block which originally come from calibration data; here, the pruning mask $\bar{m}$ is fixed assuming that it is already obtained from an arbitrary pruning method. I.e., the goal is to update variables in each block to minimize the extended reconstruction errors. We solve this problem iteratively using the standard gradient-based method. Notably a similar approach is also proposed in the concurrent work of Guo et al. (2024), and we find in our experiments that br is extremely effective in reducing the reconstruction errors in Section 3.1. We illustrate the idea of br in Figure 2.
@@
Another way we consider to further reduce reconstruction errors is to extend the reconstruction unit from a block to multiple blocks and stitch the solutions in between by connecting via the adjacent blocks. Specifically, this means that now $g$ in (2) becomes a composite of multiple blocks, say $h$, and we ensure $h$ overlaps; more precisely, $h_{i}=g_{i}\circ g_{i-1}$ and $h_{i+1}=g_{i+1}\circ g_{i}$ for two blocks, and so on for all blocks. This way, namely cross-block reconstruction or cr (Ding et al., 2023), we can potentially bridge between subsolutions by taking into account some interaction between adjacent blocks, and hence, reduce the compounding errors. We illustrate the idea of cr in Figure 2.
To elaborate further, the difference between br and cr is that while br is about updating parameters within a block (thus it is not concerned with how to combine subsolutions), cr takes a step further and is about stitching the subsolutions; i.e., cr updates parameters within two adjacent blocks, and when it comes to reconstructing the next block, it includes the overlapping block so that it has the effect of “stitching”. This method is found to be quite effective for reducing the error, however, we find that this method can often lead to overfitting. We discuss this in detail in Section 3.2.
@@ 3.1 Reconstruction error
We first evaluate the effectiveness of the suggested techniques in reducing the reconstruction error. Here, we focus on pruning LLaMA-7B (Touvron et al., 2023) and OPT-125M (Zhang et al., 2022) to unstructured $50\%$ sparsity with three pruning methods: SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024), and Magnitude (Han et al., 2015). For each pruning method, we examine four reconstruction strategies: layer-wise reconstruction (lr), block-wise reconstruction (br), block-wise reconstruction with global propagation (br+gp), and cross-block reconstruction with global propagation (br+gp+cr). Following the convention, we use $256$ calibration data randomly sampled from C4 (Raffel et al., 2020) each containing $1024$ tokens. We run the Adam optimizer for $10$ epochs (see Appendix A for details). The results are presented in Figure 3.
@@ 3.2 Generalization performance
At first, we find that the perplexity effectively decreases with br and gp; the value reduces across all test cases including different models, pruning methods, and datasets. Unexpectedly, however, the perplexity rather increases when we add cr despite the reduced reconstruction errors. We also observe a similar trend in zero-shot performance for Wanda and Magnitude pruning, with mean accuracy increasing by a large margin with br and gp but decreasing with cr. Interestingly, for SparseGPT, reconstruction techniques do not generally help zero-shot performance. We hypothesize that it is because SparseGPT already conducts fairly heavy optimization compared to other methods, and applying further reconstruction on particular calibration data may not help improve zero-shot performance since it is more sensitive to distribution shift. Furthermore, we find that such overfiting tends to occur more for LLaMA-7B than OPT-125M (see Table 2). This is possibly due to model size; i.e., given the same amount of (limited) calibration data, over-optimizing can make large models more likely to overfit and lead to poor generalization.
We can summarize our findings are as follows.
  • br and gp are found to be very effective in reducing perplexity in all cases; on the other hand, cr often leads to overfitting, especially for large models.
  • This holds true for zero-shot performance as well, with only exception of SparseGPT, for which br and gp do not help much in improving zero-shot performance; this is possibly due to the fact that SparseGPT already conducted fairly heavy optimization of remaining weights. It is also possible that adapting to downstream task is more prone to overfitting. This certainly requires more investigations.
In short, we can attempt to say without much loss of generality that “br and gp can generally help for pruning LLMs in terms of reducing perplexity”.
@@ 4 Further Exploration
We have seen that reconstruction techniques are useful but they can lead to undesirable overfitting. Here we explore potential ways to alleviate this risk. In particular, we identify that the calibration data is highly limited in two aspects: it is too little (compared to optimization variables)222 This can be especially problematic for domain-specific LLMs, e.g., healthcare (Singhal et al., 2023; Luo et al., 2022) and finance (Wu et al., 2023; Yang et al., 2023), where obtaining real-world data can be highly challenging due to privacy concerns. and does not represent the training data (as it is arbitrarily given); the former is related to the general representation-generalization complexity trade-off, and the latter is about whether the reconstruction can mimic the behavior of the original model.
To this end, we reflect on the fact that what we are dealing with is a generative (language) model, meaning that we can create calibration data that is potentially much bigger in size and closer to the original distribution. We find that this self-generation technique has recently been proposed in other contexts (Meng et al., 2022; Ye et al., 2022; Liu et al., 2023; Li et al., 2024), and thus, follow the process therein to produce high-quality text data. Using that, we perform reconstruction again, and the results are reported in Figure 4. We observe that making use of more self-generated calibration data (without unfairly violating the given setting) reduces both test error and perplexity, mitigating overfitting quite effectively.
@@ 6 Limitations
There remain several limitations in our experiments and we plan to address these in future work. First, our main experiments are limited to LLaMA-7B and OPT-125M. We intend to scale up our experiments to much larger models of up to 70B parameters and different architectures including Mixtral (Jiang et al., 2024) or Gemma (Team et al., 2024). Next, reconstruction techniques br, gp, and cr require additional memory compared to lr, although they still use much less memory compared to model-level reconstruction of solving (1) (see Appendix B for the details). We plan to introduce parameter-efficient optimization (Hu et al., 2022) to alleviate this increased memory burden.
Although the self-generation of calibration data effectively mitigates overfitting, it requires more computation for reconstruction. Finally, we find that some portions of the generated texts are far from plain English texts and thus may not serve as good calibration data (see Table 5 of Appendix C for the examples). In this regard, we believe that reducing the number of these irrelevant examples and generating only a few number of high-quality texts can be a potential way to improve performance and increase efficiency.
Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH); RS-2022-II220959/No.2022-0-00959, (part2) Few-Shot learning of Causal Inference in Vision and Language for Decision Making; RS-2024-00338140, Development of Learning and Utilization Technology to Reflect Sustainability of Generative Language Models and Up-to-Dateness over Time) and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2023-00210466, RS-2023-00265444, RS2023-0021371). Sungbin Shin was supported by Kwanjeong Educational Foundation Scholarship.
@@ Appendix B Additional Results
lr br br+gp br+gp+cr Full fine-tuning
peak memory (GB) $3.9$ $5.7$ $5.7$ $10.6$ $>100$
Table 6: Peak GPU memory for LLaMA-7B and sparseGPT. Compared to lr, reconstruction techniques incur additional GPU memory but it is quite marginal compared to fine-tuning the full model. The results are obtained with the batch size of $8$ and gradient accumulation. For full fine-tuning, the results are from Malladi et al. (2023).
Pruning block lr br br+gp br+gp+cr
Attention $32.82$ $30.15$ $29.97$ $29.64$
Feed-forward $30.69$ $29.23$ $28.89$ $28.73$
All $36.35$ $31.94$ $31.57$ $30.86$
Table 7: Effects of pruning block for different reconstruction techniques. Here, we prune either attention or feed-forward block to $50\%$ sparsity and measure the perplexity on raw-Wikitext2. Pruning only the attention block leads to worse performance compared to pruning only the feed-forward block. The results are for OPT-125m with sparseGPT.
More results on the reconstruction techniques
@@
Memory consumption of reconstruction techniques
Solving (1) directly can be memory-intensive, thus many recent work suggest divide-and-conquer such as lr and br. In the work of Frantar and Alistarh (2023), the authors show that for the 175B parameter OPT model it requires at least five A100 GPUs of 80GB, whereas by using lr it reduces down to a single A100 GPU of 80GB. In our experiments, for Llama-7B, both lr and br+gp+cr can all be done on a commodity 3090 GPU of 24GB memory; it requires more than 100GB to perform full fine-tuning of LLaMA-7B (Malladi et al., 2023). In theory, optimizing more parameters can incur more memory footprints, and thus, in the order of lr $=$ gp $<$ br $<$ cr, there will be more memory usage.
The exact amount depends on the specific model. To provide solid evidence, we ran profiling peak GPU memory for LLaMA-7B with the batch size of $8$ (see Table 6 for the results). Compared to lr, reconstruction techniques surely incur additional GPU memory, however, (i) it is quite marginal compared to fine-tuning the full model, and (ii) it could be reduced further by introducing memory reduction techniques in practice such as CPU offloading and gradient checkpointing.
Pruning attention vs. feed-forward
We also investigated the effects of only pruning attention vs. feed-forward blocks for different reconstruction techniques. Here, we conducted experiments for OPT-125m and SparseGPT by pruning either attention or feed-forward blocks to $50\%$ sparsity and measuring the perplexity on raw-Wikitext2. The results are provided in Table 7. We first observe that pruning both attention and feed-forward yields the largest performance drop. Also, we find that pruning only the attention block leads to worse performance compared to pruning only the feed-forward block, which is consistent with the findings in the previous work (Namburi et al., 2023). Interestingly, we find that reconstruction techniques can be more effective for cases with poor performance; i.e., in the order of pruning all blocks > pruning attention > pruning feed-forward, br, gp, cr reconstruction techniques yield more reduction in perplexity (which is good by itself).